OpenStudy (anonymous):

Help check my answer? im a bit confused. Factor x2 - 7x + 8. (4 points) (x + 8)(x - 1) Prime (x - 8)(x - 1) (x + 8)(x + 1)

OpenStudy (anonymous):

i guessed A

OpenStudy (anonymous):

are you in flvs

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Florida virtual school

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'm not 100% sure but I would go with A

OpenStudy (anonymous):

okay

OpenStudy (freckles):

Factor x2 - 7x + 8. @kinserkara can you tell me what two factors of 8 have product 8 and sum -7?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

its prime!

OpenStudy (freckles):

you can't think of two such factors, right? for example here are the possibles: product=8 and sum=-7 8(1) but 8+1=? -8(-1) but -8-1=? 2(4) but 2+4=? (-2)(-4) but -2-4=?

OpenStudy (anonymous):

bc there are no factors that go into -7 and 8

OpenStudy (freckles):

right none of those sums are -7

OpenStudy (anonymous):

(x-8)(x+1) since there is an x squared, then you will have an x in each parenthesis. Now you need to find 2 numbers that will add to -7 AND multiply to -8. these numbers will each go in the parenthesis. In this problem, the two numbers are -8 and 1

OpenStudy (anonymous):

thank you all